The difference between the circumference and radius of a circle is 37 cm. find the area of the circle.
Answer by amarjeeth123(569)

Let the radius be r.
Circumference of the circle=2*pi*r
2*pi*r-r=37
6.283185r-r=37
5.2831r=37
r=7.0033 cm